Find the equation.
The line with the x-intercept 5 and y-intercept -7.

Answer by tutorcecilia(2152)

The line with the x-intercept 5 and y-intercept -7
Use the slope-intercept equation of the line:
y=mx+b
y-intercept = b = (0, -7)
x-intercept = (5, 0)
Plugging in the values of (5, 0):
0=(m)(5)-7 [solve for the m=value]
0=5m-7
7=5m
7/5=m
.
y=mx+b [plug-in the slope and the y-intercept (b)]
y=7/5bx - 7